PLEASE WHEN ANSWERING THIS PROBLEM PROVIDE SOLUTIONS AND STEPS. (Especially if using excel). I am trying to understand the material. I would really appreciate if you could give me a step by step answer. Thank you in advance....

You are about to start to consider a batch of new capital budgeting projects. Before you begin, you need to estimate your company’s Weighted-Average-Cost-of-Capital (WACC). The firm operates in the 35% marginal tax bracket. There are foursets of liability holders on the balance sheet. Calculate the WACC including all four classes of liabilities.

A.There are 7,630,000 shares of common stock outstanding. These are trading at $43.56 per share. You have decided to use the Gordon Growth Model to estimate the return

required by your firm’s shareholders. In the most recent annual report the earnings per share (EPS) were $3.25. You feel that it is reasonable to assume that earnings will grow

at about 1.20% into the foreseeable future.

B. There is an issue of 825,000 shares of preferred stock. These securities promise an(perpetual) annual dividend of $7.15. The shares are currently trading for $97.36.

C.There is an issue of 145,000 coupon bonds, which have a face value of $1,000, pay 4.75% (annual) coupons, and mature in 7 years. The securities are currently trading for $930.22.

D.The firm also has 50 short-term commercial paper notes outstanding that have a face value of $1,000,000, and mature in 49 days. These are currently selling for $995,392.37.
Answer #1

Sol. To calculate the WACC for all the four liabilities,we need the return for each liability based on their current return and current trading price.

1) for common stock:

total value= quantity * price=$332,362,800

return=EPS/ price=3.25/43.56=7.46%

2) for preferred stocks:

total value=825000*97.36=$80,322,000

return=7.15/97.36=7.34%

3)for coupon bonds:

total value=145000*930.22=$134,881,900

return=4.75%*1000/930.22=5.11%

4)FOR short term commercial paper:

total value=$995,932.37

annual return= (1000000-995392.37)*(365/49)/995392.37=3.45%

total value=$548,562,632.4

WACC= return on stocks*fraction of total value+ return on debt(1-tax rate)*fraction of total value

= 7.46*0.61 + 7.34*0.15 + (1-0.35)(5.11*0.25 + 3.45*0.0018)=6.48%